Prepare a pro forma income statment for 2016. Assume that the sales increase to $87 million as a result of the plant expansion. Also assume that the cost of sales and selling and administrative expense rations (as a percentage of sales) remian constant. Finally, assume that interest expense and the firm's tax rate remain the same in 2016.TABLE 160.1 Anderson Furniture Companys Financi urniture Companys Financial Data (in Thousands of Dollars) Industry Average
Answer #1

Anderson Furniture Company (in thousands of dollars)

Proforma income statement for 2016:

$ in thousands Percentage
Net sales 87,000 100%
Less: Cost of sales (70,470) (81%)
Gross profit 16,530 19%
Less: Selling and Administration expenses (8,700) (10%)
Earnings before interest and taxes 7,830 9%
Less: interest (2,100) (2.4%)
Earnings before taxes 5,730 6.6%
Less: Tax @ 45.16% (2,588) (3%)
Earnings after tax 3,142 3.6%

Workings:

Cost of sales = 87,000 × 81% = $ 70,470

Gross profit = Net sales - cost of sales = 87,000 - 70,470 = 16,530

Selling and Administration expenses = 87,000 × 10% = 8,700

EBIT = gross profit - selling and administration expenses = 16,530 - 8,700 = 7,830

Percentage of interest = Interest / Net sales = 2,100/87,000 = 2.4%

Tax = 45.16% (EBT) = 45.16% (5,730) = 2,588

Percentage of tax = 2,588/87,000 = 3% (approximately)

Earnings after tax = EBT - tax = 5,730 - 2,588 = 3,142

Percentage of EAT = 3,142/87,000 = 3.6%